the average Mark of candidates in an aptitude test was 128.5 with a standard deviation of 8 points to the three scores extracted from the test are 148 102 and 152 what is the average of the extractor scores that are extreme values
1
Expert's answer
2020-06-02T18:27:18-0400

"\text{According to Z- score }\\
\text{any data value with a z-score less}\\
\text{ than -2 or greater than 2 as an outlier}\\
Z=\frac{x- \mu}{\sigma}\\
Z_{148}=\frac{148- 128.5}{8.2}=2.38\\
Z_{102}=\frac{102- 128.5}{8.2}=-3.23\\
Z_{152}=\frac{152- 128.5}{8.2}=2.86\\
\text{ so all of them are outliers}\\
\text{the average of the extracted scores that }\\
\text{are extreme values (outliers)}
\\=\frac{148+102+152}{3}=134"